OpenStudy (anonymous):

solve the indicated letter b=4c, for c

OpenStudy (anonymous):

c=b/4

OpenStudy (anonymous):

is tht c=b*4 or you mean c=b/4 as in divide

OpenStudy (anonymous):

c=b/4 as in divide